OpenStudy (anonymous):

y=(4-x^2)(x^3-x+1) find y' both by multiplication and product rule

OpenStudy (anonymous):

really? myininaya, which one do you want?

myininaya (myininaya):

\[\ln(y)=\ln[(4-x^2)(x^3-x+1)]\] i want the easy way lol

OpenStudy (anonymous):

give me a break

myininaya (myininaya):

ok you choose what way you want i will do the other

OpenStudy (anonymous):

\[y=-x^5+5 x^3-x^2-4 x+4\] \[y'=-5x^4+15x^2-2x-4\]

myininaya (myininaya):

\[y'=(4-x^2)'(x^3-x+1)+(4-x^2)(x^3-x+1)' \] \[=(0-2x)(x^3-x+1)+(4-x^2)(3x^2-1+0) \]

OpenStudy (anonymous):

hmmm i wonder if they are the same...

myininaya (myininaya):

\[=-2x(x^3-x+1)+(4-x^2)(3x^2-1)\]

myininaya (myininaya):

\[=-5x^4+15x^2-2x-4\]

myininaya (myininaya):

yes they are

myininaya (myininaya):

\[-2x^4+2x^2-2x+12x^2-4-3x^4+x^2=x^4(-2-3)+x^2(2+12+1)-2x-4\]

myininaya (myininaya):

\[=-5x^4+15x^2-2x-4\]